# Learning to Reconstruct Quirky Tracks

### v1

@@ VII Conclusions
Recent advances in ML-based reconstruction allow for new inroads into the previously inaccessible territory. In this work we have shown that graph neural network tracking can be trained to reconstruct quirk tracks with macroscopic oscillation lengths in the problematic centimeter range. We have generated quirk pseudodata across the $m_{Q},\Lambda$ parameter space in a simplified model of the ATLAS detector and trained the Exa.TrkX machine learning algorithm to reconstruct quirks. When trained on only SM tracks, the algorithm naturally fails to reconstruct tracks, but when instead trained on quirks the algorithm reaches comparatively large reconstruction efficiencies, $\sim 40-90\%$. We have also tested the robustness of the algorithm, by including background tracks from the associated jet in the signal process and minimum bias events arising from pileup and found that these additional backgrounds do not significantly degrade our reconstruction performance. The agnostic nature of the ML pipeline to the specific features of these tracks suggests a potentially broader capability to reconstruct non-helical tracks in a model-independent way.
@@
VIII Acknowledgements
DW is funded by the DOE Office of Science. Y.Fang and Q.Sha are supported partially by Institute of High Energy Physics, Chinese Academy of Sciences under the innovative project on sciences and technologies with no.E3545BU210. DM is supported by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429). The work of MF was supported in part by U.S. National Science Foundation Grants PHY-2111427 and PHY-2210283, as well as by NSF Graduate Research Fellowship Award No. DGE-1839285. The authors are grateful to Simon Knapen for useful discussions and for sharing the quirk trajectory code from  Knapen et al. (2017). We are also grateful to Markus Elsing and Xiangyang Ju for their comments and expertise.
@@ Appendix A SM efficiency
5000 29.8% - - -

### v2

@@ VIII Conclusions
Recent advances in ML-based reconstruction allow for new inroads into the previously inaccessible territory. In this work we have shown that graph neural network tracking can be trained to reconstruct quirk tracks with macroscopic oscillation lengths in the problematic centimeter range. We have generated quirk pseudodata across the $m_{Q},\Lambda$ parameter space in a simplified model of the ATLAS detector and trained the Exa.TrkX machine learning algorithm to reconstruct quirks. When trained on only SM tracks, the algorithm naturally fails to reconstruct tracks, but when instead trained on quirks the algorithm reaches comparatively large reconstruction efficiencies, $\sim 40-90\%$. We have also tested the robustness of the algorithm, by including background tracks from the associated jet in the signal process and found that these additional backgrounds do not significantly degrade our reconstruction performance or the sensitivity to discovering quirks. Effects from pile-up remain to be studied rigorously in future work, but existing implementations of GNN reconstruction in high pile-up (HL-LHC-like) conditions are performant with comparable sensitivity to traditional techniques Ju et al. (2021); Caillou et al. (2022). Although quirk trajectories are clearly more topologically complex than an SM hadron, we have no reason to suspect that their reconstruction is particularly sensitive to pile-up and noise effects, and remain optimistic that future studies in this direction will maintain the high efficiency demonstrated in this work. The agnostic nature of the ML pipeline to the specific features of these tracks suggests a potentially broader capability to reconstruct non-helical tracks in a model-independent way.
@@
IX Acknowledgements
DW is funded by the DOE Office of Science. Y.Fang and Q.Sha are supported partially by Institute of High Energy Physics, Chinese Academy of Sciences under the innovative project on sciences and technologies with no.E3545BU210. DM is supported by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429). The work of MF was supported in part by U.S. National Science Foundation Grants PHY-2111427 and PHY-2210283, as well as by NSF Graduate Research Fellowship Award No. DGE-1839285. The authors are grateful to Simon Knapen and Michele Papucci for useful discussions and for sharing the quirk trajectory code from  Knapen et al. (2017). We are also grateful to Markus Elsing and Xiangyang Ju for their comments and expertise.
@@ Appendix A SM efficiency
5000 29.8% 24.6% - -
Table 3: Score cut for different points
$m_{Q}$ (GeV) $\Lambda$(eV) $\bar{\gamma}$ $\sigma_{\gamma}$ $d$ [cm] Efficiency Score cut
100 100 4.4 3.4 670 91.0% 0.9
500 3.7 2.7 21.6 82.8% 0.9
1000 3.1 2.05 4.2 77.8% 0.9
2000 3.0 2.0 1.0 60.4% 0.6
3000 3.0 1.9 0.44 35.6% 0.5
4000 2.9 1.8 0.24 34.5% 0.5
5000 2.9 1.7 0.15 24.2% 0.5
500 100 1.9 0.7 896 92.0% 0.9
500 1.8 0.6 31 79.0% 0.9
1000 1.7 0.6 7.3 64.3% 0.9
2000 1.7 0.5 1.7 60.9% 0.6
3000 1.7 0.5 0.8 59.6% 0.6
4000 1.6 0.5 0.4 42.6% 0.5
5000 1.6 0.5 0.2 39.2% 0.5
1000 100 1.5 0.3 950 94.6% 0.9
500 1.4 0.3 32 63.6% 0.9
1000 1.4 0.3 7.6 62.7% 0.6
2000 1.4 0.3 1.8 69.2% 0.7
3000 1.3 0.2 0.8 54.1% 0.6
4000 1.3 0.2 0.4 59.7% 0.6
5000 100 1.04 0.03 420 84.8% 0.9
500 1.03 0.02 11 69.8% 0.9
1000 1.03 0.02 2.7 69.5% 0.6
2000 1.03 0.02 0.7 49.2% 0.5
3000 1.03 0.02 0.3 36.4% 0.5
4000 1.03 0.02 0.2 34.6% 0.3